OpenStudy (anonymous):

The sum of five consecutive integers is 75. What is the third number? 13 14 15 16

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

Solve the following for x: x - 2 + x - 1 + x + x + 1 + x + 2 =75 x = 15

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

wrong

OpenStudy (anonymous):

explain how

OpenStudy (anonymous):

jk

OpenStudy (anonymous):

it is 15 i asked people he is right give him a medal

OpenStudy (anonymous):

13 + 14 + 15 + 16 + 17 = 75

OpenStudy (anonymous):

how old are you @robtobey

OpenStudy (anonymous):

hes 80??

OpenStudy (anonymous):

lol he look like 15

OpenStudy (anonymous):

80 next August. The photo used for openoffice is my 6th grade graduation photo taken in 1946.

OpenStudy (anonymous):

2 = 2t – 2 @robtobey

OpenStudy (anonymous):

t = 2

OpenStudy (anonymous):

3n + 1 = 10

OpenStudy (anonymous):

Sorry, openoffice should have been written, openstudy.

OpenStudy (anonymous):

n = 3

OpenStudy (anonymous):

10 – 2f = 6 @robtobey

OpenStudy (anonymous):

f = 4

OpenStudy (anonymous):

2b + 4 = 10

OpenStudy (anonymous):

After 1AM here, Got to bail out.

OpenStudy (anonymous):

2b + 4 = 10

OpenStudy (anonymous):

@robtobey